# Predicting the duration of traffic incidents for Sydney greater metropolitan area using machine learning methods

### v1

@@ Abstract
This research presents a comprehensive approach to predicting the duration of traffic incidents and classifying them as short-term or long-term across the Sydney Metropolitan Area. Leveraging a dataset that encompasses detailed records of traffic incidents, road network characteristics, and socio-economic indicators, we train and evaluate a variety of advanced machine learning models including Gradient Boosted Decision Trees (GBDT), Random Forest, LightGBM, and XGBoost. The models are assessed using Root Mean Square Error (RMSE) for regression tasks and F1 score for classification tasks.
@@
The proposed methodology not only achieves high predictive accuracy but also provides stakeholders with vital insights into factors contributing to incident durations. These insights enable more informed decision-making for traffic management and response strategies. The code is available by the link: https://github.com/Future-Mobility-Lab/SydneyIncidents
@@ 1 Introduction
Efficient prediction and management of road traffic incidents are crucial in minimizing congestion and enhancing safety, particularly in densely populated areas. The precision of these predictions profoundly impacts emergency response times and overall traffic management strategies. Central to these predictions is the estimation of ’incident duration’. The term ’incident duration’ can refer to different time spans. Some studies define it as the interval from the onset of an incident to the clearance of the scene from the main road, while others consider it the period until the restoration of normal traffic conditions. The latter definition often requires more detailed analyses of traffic conditions and traveller behaviour. Both durations are influenced by factors such as the nature of the incident, the effectiveness of the emergency responses, and the prevailing road and traffic conditions.
@@
The practical applications of accurate incident duration predictions are extensive. For emergency response units, including ambulance and fire services, these predictions facilitate strategic planning concerning the placement and readiness of teams. Traffic management authorities also use these insights to optimize the number and location of response units, effectively reducing incident durations. Furthermore, integrating real-time traffic simulation or traffic predictive systems with incident duration models can revolutionize traffic management strategies. These integrations allow for dynamic adjustments of traffic signals and route advisories based on ongoing incidents, showcasing the transformative potential of advanced predictive analytics in urban traffic management. Historical incident data plays a pivotal role in developing predictive models. While some studies focus on exact duration predictions, others classify incident durations into categories such as short-term or long-term, aiding in rapid and effective response strategies. Various models, ranging from simple regression techniques to sophisticated machine learning algorithms like Gradient Boosted Decision Trees, Random Forest, and XGBoost, are employed to estimate incident durations. These models not only predict the time needed to resolve incidents but also assist in resource and response planning by classifying incident durations appropriately.
@@
The effectiveness of these models largely depends on the granularity and relevance of the data used. Recent compilations of detailed road incident data, particularly for the Sydney metropolitan area, offer a unique opportunity to refine these models. This study represents one of the pioneering efforts to apply state-of-the-art machine learning techniques to such comprehensive datasets, aiming to set new benchmarks in the accuracy and reliability of incident duration predictions. By focusing on Local Government Area (LGA) data within the Sydney Metropolitan Area, this study explores the relationship between incident duration and local characteristics. Leveraging this detailed dataset, we aim to provide insights that enhance the development of predictive models, thereby improving traffic management strategies across varied urban landscapes.
@@
This local focus is crucial as it acknowledges the specific traffic distribution, demand, and road network features of Sydney, addressing a critical gap in existing research and capturing the unique characteristics of road traffic incidents that can vary significantly from one region to another.
In this study, we undertake a comprehensive examination of the factors influencing the duration of traffic incidents within the Sydney Metropolitan Area. Leveraging a detailed dataset that encompasses traffic incidents, road characteristics, and socio-economic indicators, we employ advanced machine learning models, including Gradient Boosted Decision Trees, Random Forest, LightGBM, and XGBoost. Our primary objective is to predict the duration of traffic incidents and classify them as either short-term or long-term. Our analyses reveal that models such as XGBoost and LightGBM notably outperform traditional approaches, with XGBoost achieving the lowest RMSE for predicting incident durations and the highest F1 score for classification tasks. We also conduct an extensive feature importance analysis using methods like tree split counts and SHAP values, highlighting the significance of factors such as the number of affected lanes, traffic volume, and types of vehicles involved.
@@
Incident duration prediction includes several critical timelines: the time required for emergency services or incident authorities to arrive at the scene, the time taken to clear the incident from the roadway, and the time until traffic flow returns to pre-incident conditions.
@@ 2 Related Works
In the last two decades, various approaches have been developed, involving different machine learning strategies and data parameters.
@@
In contrast, efforts by Chang and Chang [7] to use a classification tree model were less successful. While the classification accuracy for short incident durations was very high, the model struggled to similarly classify medium and long duration incidents. Dimitriou and Vlahogianni [7], analysing the same data set, concluded that increasing the number of input variables does not necessarily improve the model performance.
@@
Critically, most of the existing literature on traffic incident duration predictions overlooks the role of localized traffic characteristics, with only 20% of reviewed studies employing European traffic data, according to Li et al. [9]. Moreover, these studies have not yet achieved sufficient accuracy levels to offer real operational benefits.
Considering the varied and complex nature of traffic incidents, several studies have opted to concentrate their research on certain types of incidents, such as vehicle collisions or mechanical failures [9]. This focused methodology permits an in-depth examination of the distinct attributes inherent to these incidents, recognizing that a universal approach might not effectively encompass the intricacies of different incident categories. The rationale behind specializing in particular incident types is to bolster the predictive accuracy of the models and enhance their effectiveness in situations where general models may not perform as well. Often, the scarcity of detailed data on specific types of incidents can result in predictive inaccuracies concerning traffic conditions. Moreover, several investigations have shed light on the environmental consequences associated with the traffic incidents duration, uncovering how the duration of such events interacts with environmental effects and traffic dynamics. These studies deliver essential insights into emissions, energy consumption, and other critical factors, yielding vital directions for policymakers and researchers focused on minimizing the environmental impact of traffic incidents and promoting the sustainability of transportation networks. For a more in-depth examination of incident duration and its ramifications, recent comprehensive reviews, such as those found in sources X and Y, provide extensive analysis and discussion.
@@ 3 Case Study: Incident data and Metadata of Sydney Metropolitan Area
Figure 1: Heatmap of traffic incidents with marked centroids
@@ 31. SA4 NAME21 Statistical Area Level-4, Zone name "City Region A", "City Region B", …
The metadata (see Table 3) encompasses a wide range of urban and socio-economic indicators:
@@ 4 Data exploration
When exploring and comparing the distributions of incident durations across various categorical variables, an intuitive approach is to utilize Wasserstein distance. The Wasserstein distance (also known as Earth Mover’s Distance) is a measure of the distance between two probability distributions. It provides a quantitative way to compare the shapes and positions of distributions, which is particularly useful in highlighting differences in distributions of incident durations across different categorical groups.
The Wasserstein distance, also known as the Earth Mover’s Distance, arises from optimal transport theory, which is concerned with finding the most efficient way to reshape one probability distribution into another. This concept can be visualized as the minimum "work" needed to transport a pile of dirt configured in one shape to match another shape.
@@
Figure 11 shows the heatmap of average Wasserstein distances for Number of Vehicles Involved. There is a low amount of difference observed for incident duration patterns across small numbers of vehicles involved. Significant deviation of patterns starts when at least 5 vehicles involved in accident, with extreme variation in duration patterns for incidents involving 7 to 9 vehicles.
@@ 5 Methodology
    Feature Importance based on Decision Tree splits: when using Decision Trees we can calculate the importance of features by measuring how much each feature decreases the impurity of a split (for classification) or the variation (for regression). Features that are selected more often for splitting at the top of the trees are considered more important.
@@ 6 Results
The tables provided represent confusion matrices at different time thresholds for predicting if a traffic incident will be short-term or long-term. Each matrix categorizes incidents into four parts based on Actual and Predicted durations:
True Positive (TP): Traffic incidents correctly predicted as short-term (upper-left quadrant). False Negative (FN): Traffic incidents wrongly predicted as long-term but were actually short-term (upper-right quadrant). False Positive (FP): Traffic incidents wrongly predicted as short-term but were actually long-term (lower-left quadrant). True Negative (TN): Traffic incidents correctly predicted as long-term (lower-right quadrant).
@@
In cases of traffic incident duration classification performance (see Figure 15), XGBoost and LightGBM show the highest Accuracy and F1-score (0.67 and 0.62 correspondingly). The same performance is observed for regression task (see Figure 16, RMSE is 33.7 for XGBoost).
Figure 15: Comparative performance of classification models for selected duration threshold of 30 minutes
Figure 16: Comparison of model RMSE values.
@@ 8 Conclusion
The feature importance analysis employed two distinct approaches: feature importance based on tree splits and SHAP values. The XGBoost model identified key features influencing predictions as the number of affected lanes, the percentage of dwellings with four or more motor vehicles (Feature 38_ PD4MV), the scale of the incident, traffic volume, and motorway length. These features were frequently used in decision-making splits within the XGBoost trees.
@@
The harmonious results from both the tree splits and SHAP analysis validate the significance of traffic volume, the number of affected lanes, and related features in predicting traffic incident durations. This study underscores the importance of feature selection and threshold setting in improving model accuracy and achieving balanced prediction performance for traffic incident duration classification.
In conclusion, the findings demonstrate the effectiveness of machine learning models in predicting traffic incident durations, with XGBoost and LightGBM offering robust performance for classification. The feature importance analysis provides valuable insights, highlighting the critical factors influencing incident duration predictions. These insights can guide future efforts in developing more accurate and balanced traffic incident management systems.

### v2

@@ Abstract
This research presents a comprehensive machine learning approach to predicting the duration of traffic incidents, classifying them as short-term or long-term, and understanding what are the factors that affect the duration the most. Our modelling methodology is using a dataset from the Sydney Metropolitan Area that includes detailed records of traffic incidents, road network characteristics, and socio-economic indicators, we train and evaluate a variety of advanced machine learning models including Gradient Boosted Decision Trees (GBDT), Random Forest, LightGBM, and XGBoost. The models are assessed using Root Mean Square Error (RMSE) for regression tasks and F1 score for classification tasks.
@@
The proposed methodology not only achieves high predictive accuracy but also provides stakeholders with vital insights into local area factors contributing to incident durations. These insights enable more informed decision-making for traffic management and response strategies. The code is available by the link: https://github.com/Future-Mobility-Lab/SydneyIncidents
@@ 1 Introduction
Efficient prediction and management of road traffic incidents are crucial in minimizing congestion and enhancing safety, particularly in densely populated areas. The precision of these predictions profoundly impacts emergency response times and overall traffic management strategies. Central to these predictions is the estimation of ’incident duration’. The term ’incident duration’ can refer to different time spans. Some studies define it as the interval from the onset of an incident to the clearance of the scene from the main road by the response team, while others consider it the period until the restoration of normal traffic conditions. The latter definition often requires more detailed analyses of traffic conditions and traveller behaviour. Both durations are influenced by factors such as the nature of the incident, the effectiveness of the emergency responses, and the prevailing road and traffic conditions.
@@
The practical applications of accurate incident duration predictions are extensive. For emergency response units, including ambulance and fire services, these predictions facilitate strategic planning concerning the placement and readiness of teams. Traffic management authorities also use these insights to optimize the number and location of response units, effectively reducing incident durations. Furthermore, integrating real-time traffic simulation or traffic predictive systems with incident duration models can revolutionize traffic management strategies. These integrations allow for dynamic adjustments of traffic signals and route advisories based on ongoing incidents, showcasing the transformative potential of advanced predictive analytics in urban traffic management. Historical incident data plays a pivotal role in developing predictive models. While some studies focus on exact duration predictions, others classify incident durations into categories such as short-term or long-term, aiding in rapid and effective response strategies. Various models, ranging from simple regression techniques to sophisticated machine learning algorithms like Gradient Boosted Decision Trees, Random Forest, and XGBoost, are employed to estimate incident durations. These models not only predict the time needed to resolve incidents but also assist in resource and response planning by classifying incident by durations appropriately.
@@
The effectiveness of these models largely depends on the granularity and relevance of the data used. Recent compilations of detailed road incident data, particularly for the Sydney metropolitan area, offer a unique opportunity to refine these models. This study represents one of the pioneering efforts to apply state-of-the-art machine learning techniques to such comprehensive datasets, aiming to set new benchmarks in the accuracy and reliability of incident duration predictions. By focusing on Local Government Area (LGA) data within the Sydney Metropolitan Area, this study explores the relationship between incident duration and local road parameters, economic and travel behaviour characteristics. Leveraging this detailed dataset, we aim to provide insights that enhance the development of predictive models, thereby improving traffic management strategies with the consideration of these variables.
@@
This local focus is crucial as it acknowledges the specific traffic distribution, demand, and road network features of Sydney, addressing a critical gap in existing research and capturing the unique characteristics of road traffic incidents that can vary significantly from one local area to another.
@@
In this study, we undertake a comprehensive examination of the factors influencing the duration of traffic incidents within the Sydney Metropolitan Area. Our primary objective is to predict the duration of traffic incidents and classify them as either short-term or long-term. Our analyses reveal that models such as XGBoost and LightGBM notably outperform traditional approaches, with XGBoost achieving the lowest RMSE for predicting incident durations and the highest F1 score for classification tasks. We also conduct an extensive feature importance analysis using methods like tree split counts and SHAP values, highlighting the significance of factors such as the number of affected lanes, traffic volume, and types of vehicles involved.
@@ 2 Related Works
In the last two decades, various approaches have been developed, involving different statistical models, machine learning approaches and data sources.
@@
In contrast, efforts by Chang and Chang [10] to use a classification tree model were less successful. While the classification accuracy for short incident durations was very high, the model struggled to similarly classify medium and long duration incidents. Recent works on traffic incident duration prediction utilise more advanced tree-based models [17, 20, 6].
@@
Most of the existing literature on traffic incident duration predictions overlooks the role of location-specific traffic characteristics (like proximity of various amenities), with only 20% of reviewed studies using data on traffic condition [12]. However, majority of these studies have not yet achieved sufficient accuracy levels to offer real operational benefits, which underlines the role of statistical analysis. Feature importance estimation methods across traffic incident duration prediction studies include the use of XGBoost tree split analysis [8], combination of tree-based prediction models with SHAP [28, 27]. Availability of data on traffic flow can also support or even enhance the incident duration prediction accuracy [20, 6].
Despite the use of detailed incident report datasets, majority of studies do not combine multiple types of data sources (i.e., social and economic data, event data, points-of-interest data, weather data, etc) [14] and rely primarily on incident report data. The combination of multiple types of features allows to perform more robust feature importance estimation, where incident-specific and indirectly related variables are compared against each other.
Considering the varied and complex nature of traffic incidents, several studies concentrate their research on certain types of incidents, such as vehicle collisions or mechanical failures and different aspects of incident timeline (clearance time, response time, total time of the incident) [12]. This focused methodology permits an in-depth analysis of the distinct attributes inherent to these categories of incidents or duration intervals, recognizing that a universal approach might not effectively grasp the aspects of different incident categories. The rationale behind specializing in particular incident types is to increase the predictive accuracy of the models and enhance their effectiveness in situations where general models may not perform as well. On the contrary, the scarcity of detailed data on specific types of incidents can result in low prediction accuracy making the general analysis preferable.
@@ 3 Case Study: Incident data and Metadata of Sydney Metropolitan Area
Figure 1: Heatmap of traffic incidents with marked centroids. Source: OpenStreetMap
@@ 31. SA4 NAME21 Statistical Area Level-4, Zone name "City Region A", "City Region B", …
The metadata (see Table 3) includes a wide range of urban and socio-economic indicators:
@@ 24 HA Hospital/Medical Area (km2) 49 ANP_FH Average Number Persons in a Family Household
In total, the merged data set, where each incident data associated with corresponding socio-economic and road network data contains 82 variables for each of 85,612 incident records.
@@ 4 Data exploration
When exploring and comparing the distributions of incident durations across various categorical variables, an intuitive approach is to utilize Wasserstein distance estimation. The Wasserstein distance (also known as Earth Mover’s Distance) is a measure of the distance between two probability distributions [4, 25]. It provides a quantitative way to compare the shapes and positions of distributions, which is particularly useful in highlighting differences in distributions of incident durations across different categorical groups.
The Wasserstein distance arises from optimal transport theory, which is concerned with finding the most efficient way to reshape one probability distribution into another. This concept can be visualized as the minimum "work" needed to transport a pile of dirt configured in one shape to match another shape.
@@
Figure 11 shows the heatmap of average Wasserstein distances for Number of Vehicles Involved. There is a low amount of difference observed for incident duration patterns across small numbers of vehicles involved. Significant deviation of patterns starts when at least 5 vehicles involved in incident, with extreme variation in duration patterns for incidents involving 7 to 9 vehicles.
@@ 5 Methodology
    Feature Importance based on Decision Tree splits: when using Decision Trees we can calculate the importance of features by measuring how much each feature decreases the impurity of a split (for classification) or the variation (for regression) [26, 30]. Features that are selected more often for splitting at the top of the trees are considered more important.
@@ 6 Results
For the classification and regression scenarios we report the score of 5-fold cross-validation, where 80% of data used for training each model and the rest 20% of the data used for prediction. The data is randomly shuffled before the modelling, allowing to focus on incident characteristics rather than a timeline of incident records, providing a robust estimate of model generalizability.
The figures provided represent accuracy and F1-scores at different time thresholds for predicting if a traffic incident will be short-term or long-term. Each prediction step categorizes incidents into four parts based on Actual and Predicted durations:
True Positive (TP): Traffic incidents correctly predicted as short-term (upper-left quadrant). False Negative (FN): Traffic incidents wrongly predicted as long-term but were actually short-term (upper-right quadrant). False Positive (FP): Traffic incidents wrongly predicted as short-term but were actually long-term (lower-left quadrant). True Negative (TN): Traffic incidents correctly predicted as long-term (lower-right quadrant). Precision is the ratio of correctly predicted positive observations to the total predicted positive observations. For example, precision would be the ratio of correctly identified short-term incidents to the total number of incidents classified as short-term by the model. Recall is the ratio of correctly predicted positive observations to all the observations in the actual positive class. For example, recall would be the ratio of correctly identified short-term incidents to the total number of actual short-term incidents, high recall would also mean that the model correctly identifies most of the actual short-term incidents.
Accuracy is the ratio of correctly predicted instances to the total instances. It is a common metric for evaluating classification models:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$
The F1-score is the harmonic mean of precision and recall and is often preferred over accuracy, especially in situations where the class distribution is imbalanced. It is used to balance the trade-off between precision and recall, providing a single metric to evaluate the performance of a classification model:
$F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
Or in terms of the basic four elements of the confusion matrix:
$F_{1}=2\cdot\frac{TP}{2TP+FP+FN}$
@@
In cases of traffic incident duration classification performance (see Figure 15(a)), XGBoost and LightGBM show the highest Accuracy and F1-score (0.67 and 0.62 correspondingly). The same performance is observed for regression task (see Figure 15(b), RMSE is 33.7 for XGBoost).
(a) Comparative performance of classification models for selected duration threshold of 30 minutes.
(b) Comparison of model RMSE values.
Figure 15: Comparative performance of models: (a) Classification metrics for a 30-minute duration threshold; (b) Model RMSE values.
@@ 8 Conclusion
The feature importance analysis relied on two distinct approaches: feature importance based on tree splits and SHAP values. The XGBoost model identified key features influencing predictions as the number of affected lanes, the percentage of dwellings with four or more motor vehicles (Feature 38_ PD4MV), the scale of the incident, traffic volume, and motorway length. These features were frequently used in decision-making splits within the XGBoost trees.
@@
The results from both the tree splits and SHAP analysis validate the significance of traffic volume, the number of affected lanes, and related features in predicting traffic incident durations. This study underscores the importance of feature selection and threshold setting in improving model accuracy and achieving balanced prediction performance for traffic incident duration classification.
Future research can utilize the same data but focus on spatial-temporal analysis [16] (including hotspot detection and cluster analysis) or survival models [23].
In conclusion, the findings demonstrate the effectiveness of machine learning models in predicting traffic incident durations, with XGBoost and LightGBM offering robust performance for classification. The feature importance analysis provides valuable insights into the feature importance, highlighting the critical local area factors influencing incident duration predictions. These insights can guide future efforts in developing more accurate and balanced traffic incident management systems.
Acknowledgements
This research is supported by the Australia Research Council (ARC), under scheme of the Linkage Projects (Grant ID: LP180100114).